solve for T using the rules of logarithms, round answer to three decimal places:
6^T=5^T+1

Answer by Alan3354(69443)

If the exponent of 5 is t+1,
t*log(6) = (t+1)*log(5) = tlog(5) + log(5)
tlog(6) - t*log(5) = log(5)
t = log(5)/(log(6) - log(5)) = log(5)/log(1.2)
t =~ 8.827